LBTI - Amali is the main driver. #1 goal is to verify LBTI AO is delivering what is needed for Science run at end of May
  1. Check that system is working in main science regime mag 4, 6, 8, 10, 12
    Test updated AOS tables:
    from Enrico "From our side we are progressing on the AOS table tuning for mag
    <7.5 adjusting parameter to couple with flux both on the main AO loop
    and for the pupil check too.
    We found that we were too conservative in the main AO loop and
    blinding the camera in the pupil check....
  2. As secondary goals we have:
    - LMIRCam frames acquisition from optical loop diagnostic (daytime
    test successful on monday)
    - Collect telemetry data for SR calculation from slopes
    - Test NCPA application (if never tested, we can apply a big knonw
    aberration and verify the PSF before and after)
    - Test the new Super-Rec with 576 modes
11:24 Still in foq with high huidity. Safed and powered off the LBTI WFS. No work done tonight.

-- JenniferPower - 29 Apr 2021
Topic revision: r2 - 29 Apr 2021, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback